# Multimodal Survival Modeling in the Age of Foundation Models

### v2

@@
Multimodal Survival Modeling in the Age of Foundation Models
Steven Song Center for Translational Data Science, University of Chicago, Chicago IL Department of Computer Science, University of Chicago, Chicago IL Medical Scientist Training Program, Pritzker School of Medicine, University of Chicago, Chicago IL These authors contributed equally Morgan Borjigin-Wang Center for Translational Data Science, University of Chicago, Chicago IL Google, Chicago IL These authors contributed equally Irene Madejski Center for Translational Data Science, University of Chicago, Chicago IL Department of Computer Science, University of Chicago, Chicago IL Robert L. Grossman Center for Translational Data Science, University of Chicago, Chicago IL Department of Computer Science, University of Chicago, Chicago IL Section of Biomedical Data Science, Department of Medicine, University of Chicago, Chicago IL Lead contact Correspondence: rgrossman1@uchicago.edu
SUMMARY
@@
The Cancer Genome Atlas (TCGA) has enabled novel discoveries and served as a large-scale reference through its harmonized genomics, clinical, and image data. Prior studies have trained bespoke cancer survival prediction models from unimodal or multimodal TCGA data. A modern paradigm in biomedical deep learning is the development of foundation models (FMs) to derive meaningful feature embeddings, agnostic to a specific modeling task. Biomedical text especially has seen growing development of FMs. While TCGA contains free-text data as pathology reports, these have been historically underutilized. Here, we investigate the feasibility of training classical, multimodal survival models over zero-shot embeddings extracted by FMs. We show the ease and additive effect of multimodal fusion, outperforming unimodal models. We demonstrate the benefit of including pathology report text and rigorously evaluate the effect of model-based text summarization and hallucination. Overall, we modernize survival modeling by leveraging FMs and information extraction from pathology reports.
KEYWORDS
@@
multimodal survival modeling, foundation models, text summarization, pathology reports, biomedical embeddings, TCGA
INTRODUCTION
@@
The Cancer Genome Atlas (TCGA) has been the premier cancer research resource for nearly two decades1. Throughout its history, its harmonized data2 has enabled novel discoveries through its multitudes of molecular genetic data3, histopathological images, and clinical descriptors for over 11 thousand cases across 33 cancer types. One clinical outcome that has been well studied using TCGA is cancer survival4, 5, 6.
@@
There have been numerous studies to date that have trained bespoke models to predict patient survival from the TCGA. These models have covered a wide range in the specific cancer types studied, the data modalities used as predictive features, and the model types trained. Some have leveraged histology images7, 8 or RNA sequencing (RNAseq)9, 10, 11, 12 alone, while others have explored varying degrees of multimodal integration13, 14, 15, 16, 17, 18, 19, 20, 21. Most recent papers focus on training bespoke deep-learning models9, 15, 10, 7, 11, 16, 17, 19, 20, 21, 8 and some apply simpler machine learning models14, 12, 18, however few explore the synergy of combining deep learning with simpler statistical, machine learning models13.
A modern paradigm in biomedical deep learning is the development of foundation models (FMs) to derive meaningful feature embeddings22. These FMs are trained over a large corpus of data, typically using a self supervised training objective to improve embedding generalizability to downstream modeling tasks. For example, the Universal Cell Embedding (UCE)23 model for single-cell RNAseq (scRNAseq) was trained with an adaptation of masked-language modeling (MLM)24 for gene expression data. Similarly, UNI25, an FM for digital pathology, was trained using the DINOv2 framework26 to produce robust feature embeddings.
Figure 1: Multimodal survival modeling is vastly simplified in the age of foundation models. Simple ensembles of small, linear models over zero-shot FM embeddings improve pan-cancer survival risk stratification. In our experiments, we further simplify clinical embeddings as one-hot encoded categorical features input directly to CoxPH survival models.
Biomedical text especially has seen growing development of large language model (LLM) FMs27, 28. There have been a series of adaptations of LLMs towards the biomedical domain. The early success of BioMedBERT (formerly PubMedBERT)29, a biomedical adaptation of BERT24, motivated subsequent works such as BioMedLM (formerly PubMedGPT)30 and BioMistral31, adaptations of GPT-232 and Mistral33, respectively. While these FMs were trained on text from research articles from PubMed and PubMed Central, more specific pathology-report text FMs have been developed as vision-language models (CONCH34 and MUSK35).
@@
While TCGA contains free-text data as pathology reports, these have been historically underutilized. This may in part be due to both the prior lack of interest in text data compared to the histology or sequencing data and the difficulty of working with the data format of these reports as scanned PDFs. With the current interest in applications of LLMs to biomedical domains, a recent effort used optical character recognition to extract the text from all available PDFs of the TCGA36. To the best of our knowledge, only one other work has explored the application of the TCGA reports towards survival modeling35 and differs from our approach (see Discussion).
Figure 2: Eight thousand patient cases spanning 32 cancer types with survival data, pathology reports, diagnostic slides, and gene expression.
@@
Here, we investigate the combination of classical survival models with modern FMs for predicting cancer survival using multimodal data from the TCGA, including pathology report text. Figure 1 presents a visual overview of our proposed framework. Our key findings are:
  • FM-based embeddings can be computed in low resource environments and then used as inputs to statistical or standard machine learning models. This approach does not require the extensive GPU-based computing resources needed for supervised fine-tuning of large language or foundation models, achieves comparable performance to literature35, and does not risk data leakage through fine-tuning.
  • Classical survival models predict survival using FM embeddings. We fit Cox proportional hazards (CoxPH)37 models over zero-shot embeddings extracted by FMs reduced with principal component analysis (PCA)38, 39.
  • Multimodal fusion is additive and improves survival predictions. We use a simple framework to do late-fusion of unimodal survival predictions and show that multimodal fusion outperforms unimodal models.
@@
  • Pathology report summarization improves survival predictions. We demonstrate the benefit of summarizing pathology report text for survival prediction and rigorously evaluate the effect of text hallucination.
RESULTS
Multimodal survival data across 32 cancers
We assemble a dataset of 7,982 patient cases for our study according to the inclusion flowchart presented in Figure 2. These cases span 32 of the 33 cancer types in the TCGA. We select cases that have available and valid data for survival analysis, preextracted pathology reports by Kefeli and Tatonelli36, precomputed tumor diagnostic slide embeddings by Chen et al.25, and tumor RNA-seq gene expression quantification.
For our experiments, we split the dataset into 5 cross-validation folds that are shared across all experiments. We stratify the splits by age bin, sex, race, ethnicity, observed mortality, and cancer type. Exact descriptive features of our dataset and splits are provided in Tables 7 and 8. While we stratify by cancer type and observed mortality, we do not stratify by survival duration. On average, overall survival time is comparable across splits, however per-cancer type subsets contained within each split may have too few observed mortalities for survival modeling (see Table 8).
Unimodal FM embeddings predict survival
Modality PCA Dimension
4 8 16 32 64 128 256 N/A
Demo* - - - - - - - 0.630
Canc* - - - - - - - 0.737
Expr 0.626 0.650 0.702 0.742 0.749 0.753 0.750 -
Hist 0.596 0.631 0.669 0.714 0.733 0.748 0.754 -
Text 0.565 0.610 0.690 0.725 0.745 0.751 0.752 -
Table 1: Foundation model derived embeddings predict survival. Average cross-validated C-index of CoxPH models trained over unimodal embeddings across varying PCA reductions. Demo = demographics, Canc = cancer type, Expr = BulkRNABert embedded RNA-Seq gene expression, Hist = UNI2 embedded histology, Text = BioMistral embedded pathology reports summarized by Llama-3.1-8B-Instruct. *Demographics and cancer type features were not reduced with PCA.
Modality Cancer Type (TCGA Project)
BRCA KIRC UCEC THCA LGG HNSC LUSC LUAD
Demographics 0.637 0.600 0.593 0.885 0.727 0.525 0.542 0.530
RNA-seq 0.573 0.673 0.643 0.619 0.790 0.560 0.579 0.620
Histology 0.648 0.685 0.703 0.749 0.753 0.636 0.598 0.617
Text 0.641 0.698 0.723 0.619 0.716 0.582 0.556 0.645
Multimodal 0.720 0.752 0.758 0.787 0.840 0.650 0.626 0.687
Table 2: Multimodal fusion improves pan-cancer survival prediction. Averaged C-index from 5-fold cross-validation for the top 8 most prevalent cancer types in TCGA. Text model used summarized reports. Multimodal model used fusion of all unimodal modalities. Results reported using PCA to 256 dimensions for all foundation model derived embeddings.
We fit small, linear CoxPH survival models over unimodal data and attain a peak mean cross-validated concordance index (C-index)40 of approximately 0.75 over single modalities (Table 1). For categorical demographic features or cancer type, we do no dimensionality reduction. CoxPH models fit over demographic features or cancer type alone achieve lower C-index (0.630 and 0.737, respectively) as compared to models fit over FM-derived embeddings of other biomedical data modalities.
Specifically, we fit and evaluate survival models over aggregated case-level embeddings for each modality. For FM-derived embeddings (expression, histology, and text), we tested varying PCA sizes for dimensionality reduction. We observe a general trend of increasing C-index with increasing PCA dimensions, which plateaus around PCA to 256 dimensions. Based on this observation, we report further results using PCA transformation of these embeddings to 256 dimensions.
We find that using BulkRNABert41 embeddings for tumor RNA-seq gene expression quantification, we achieve C-index = 0.750. We further find that survival modeling over UNI225 derived embeddings of tumor diagnostic histology slides results in C-index = 0.754. For pathology reports, using BioMistral31 derived embeddings of pathology reports summarized by Llama-3.1-8B-Instruct42, we attain C-index = 0.752.
With the same unimodal survival models, we evaluate their survival prediction performance across individual cancer types (Table 2). As the cancer type modality’s features would not differ for samples of the same cancer type, we do not evaluate per-cancer survival based on cancer type. We not only find that survival modeling performance of unimodal models varies across cancer types, we also observe that no single modality generally outperforms the other modalities. For example, among unimodal models, demographic-based modeling was most predictive of thyroid carcionma (THCA) survival, demographics C-index = 0.885, while the same was true for text report-based modeling of endometrial carcinoma (UCEC) survival, text C-index = 0.723.
Modality PCA Dimension
4 8 16 32 64 128 256
Expr-Hist 0.638 0.665 0.714 0.753 0.765 0.771 0.774
Expr-Text 0.636 0.671 0.724 0.762 0.774 0.779 0.778
Hist-Text 0.606 0.662 0.705 0.746 0.767 0.777 0.780
Expr-Hist-Text 0.644 0.682 0.727 0.765 0.780 0.786 0.788
Table 3: Late, multimodal fusion of unimodal, FM-based survival models improves survival prediction. For modality combinations with only FM-derived embeddings, average cross-validated C-index of multimodal CoxPH models trained over predicted risk scores from unimodal CoxPH models across varying PCA reductions. Expr = BulkRNABert embedded RNA-Seq gene expression, Hist = UNI2 embedded histology, Text = BioMistral embedded pathology reports summarized by Llama-3.1-8B-Instruct.
Modality PCA Dimension
4 8 16 32 64 128 256 N/A
Canc-Demo* - - - - - - - 0.747
Hist-Text 0.606 0.662 0.705 0.746 0.767 0.777 0.780 -
Expr-Hist-Text 0.644 0.682 0.727 0.765 0.780 0.786 0.788 -
Demo-Expr-Hist-Text 0.680 0.717 0.748 0.776 0.788 0.794 0.795 -
@@ RESULTS
Canc-Demo-Expr-Hist-Text 0.749 0.752 0.760 0.777 0.788 0.793 0.793 -
@@
Table 4: Mixing survival models of clinical features with FM-based survival models improves survival prediction. For the best combination of each number of modality combinations, average cross-validated C-index of multimodal CoxPH models trained over predicted risk scores from unimodal CoxPH models across varying PCA reductions. Demo = demographics, Canc = cancer type, Expr = BulkRNABert embedded RNA-Seq gene expression, Hist = UNI2 embedded histology, Text = BioMistral embedded pathology reports summarized by Llama-3.1-8B-Instruct. *Demographics and cancer type features were not reduced with PCA.
Multimodal fusion improves survival prediction

### v3

# Multimodal Cancer Modeling in the Age of Foundation Model Embeddings

@@
Abstract
@@
The Cancer Genome Atlas (TCGA) has enabled novel discoveries and served as a large-scale reference dataset in cancer through its harmonized genomics, clinical, and imaging data. Numerous prior studies have developed bespoke deep learning models over TCGA for tasks such as cancer survival prediction. A modern paradigm in biomedical deep learning is the development of foundation models (FMs) to derive feature embeddings agnostic to a specific modeling task. Biomedical text especially has seen growing development of FMs. While TCGA contains free-text data as pathology reports, these have been historically underutilized. Here, we investigate the ability to train classical machine learning models over multimodal, zero-shot FM embeddings of cancer data. We demonstrate the ease and additive effect of multimodal fusion, outperforming unimodal models. Further, we show the benefit of including pathology report text and rigorously evaluate the effect of model-based text summarization and hallucination. Overall, we propose an embedding-centric approach to multimodal cancer modeling.
keywords:
@@
Multimodal Cancer Data, Foundation Models, Biomedical Embeddings, TCGA, Pathology Report Summarization
Data and Code Availability
TCGA data is publicly available through the GDC API. Extracted pathology reports from Mendely, DOI: 10.17632/hyg5xkznpx.1. Precomputed embeddings of WSIs from HuggingFace: MahmoodLab/UNI2-h-features. Our anonymized source code is available from: https://github.com/StevenSong/multimodal-cancer-modeling.
Institutional Review Board (IRB)
This research does not require IRB approval.
1 Introduction
@@
The Cancer Genome Atlas (TCGA) has been the premier cancer research resource for nearly two decades (tomczak2015review). Throughout its history, its harmonized data (heath2021nci) has enabled novel discoveries through its multitudes of molecular genetic data (zhang2021uniform), histopathological images, and clinical descriptors for over 11 thousand cases across 33 cancer types. There have been numerous studies which develop deep learning methods using TCGA (sartori2025comprehensive) for various tasks, however TCGA cancer survival has been particularly well studied (abbasi2024survival; arjmand2022machine; liu2018integrated).
@@
There have been numerous studies to date that have trained bespoke models to predict patient survival from TCGA. These models have covered a wide range in the specific cancer types studied, the data modalities used as predictive features, and the model types trained. Some have leveraged histology images (wulczyn2020deep; yang2025foundation) or RNA sequencing (RNAseq) (ching2018cox; huang2020deep; qiu2020meta; nayshool2022surviveai) alone, while others have explored varying degrees of multimodal integration (chaudhary2018deep; zhan2019correlation; hao2019interpretable; ramirez2021prediction; malik2021deep; redekar2022identification; sun2023interpretable; fan2023pancancer; hao2023cancer). Most recent papers focus on training task-specific deep-learning models (ching2018cox; hao2019interpretable; huang2020deep; wulczyn2020deep; qiu2020meta; ramirez2021prediction; malik2021deep; sun2023interpretable; fan2023pancancer; hao2023cancer; yang2025foundation) and some apply simpler machine learning models (zhan2019correlation; nayshool2022surviveai; redekar2022identification), however few explore the synergy of combining deep learning with simpler statistical, machine learning models (chaudhary2018deep).
A modern paradigm in biomedical deep learning is the development of foundation models (FMs) to derive meaningful feature embeddings (bommasani2021opportunities). These FMs are typically trained over large corpora of data using self-supervision to improve generalization of embeddings to downstream tasks. Biomedical text especially has seen growing development of FM large language models (LLMs) (singhal2023large; thirunavukarasu2023large). While many LLMs have been adapted towards the biomedical domain using research text from PubMed and PubMed Central (gu2021domain; bolton2024biomedlm; labrak2024biomistral), more specific pathology-report text FMs have also been developed as vision-language models (lu2024visual; xiang2025vision).
@@
While TCGA contains free-text data as pathology reports, these have been historically underutilized. This may in part be due to the difficulty of working with the raw data format of these reports as scanned PDFs. With the current interest in applications of LLMs to biomedical domains, a recent effort by kefeli2024tcga used optical character recognition to extract the text from all available PDFs of TCGA. To the best of our knowledge, only one other work has explored the application of TCGA reports towards survival modeling (xiang2025vision) and differs from our approach (see Section 4).
@@
Here, we investigate the combination of simple machine learning models with modern FMs (Figure LABEL:fig:overview) for predicting cancer survival using multimodal data from TCGA.
Our main contributions are:
  • Quantification of pan-cancer survival prediction using rich embeddings with small models. We contemporize survival modeling for TCGA, the premier cancer research resource, using zero-shot, foundation model-derived embeddings combined with small models. Despite advanced deep learning methods, survival models over FM embeddings of single data modalities do not do substantially better than survival models over 5 tabular clinical features (see Section 3.2).
  • Simple framework for multimodal fusion of many data modalities that improves cancer survival prediction. We introduce a modular framework to do late-fusion of unimodal models, that is extensible to variable data modalities and embedding methods. Using our approach, we demonstrate that cancer data modalities are additive and non-redundant, including tabular clinical features (see Section 3.3).
@@
  • Analysis on the effect of clinical text summarization and model hallucination on cancer modeling. We develop a method for automatic pathology report summarization and propose an approach to rigorously evaluate the effect of LLM hallucinations during summarization. We demonstrate that summarizing pathology report text improves survival prediction and that LLM hallucinations do not impact survival predictions over their derived embeddings (see Sections 3.4 and 3.5).
[…]
fig:overview Refer to caption
Figure 1: Multimodal cancer modeling is vastly simplified in the age of foundation models. Simple ensembles of small, classical models over zero-shot FM embeddings improve pan-cancer survival risk stratification. Embedding methods are modular, allowing for simple experimentation and orchestration.
2 Methods
2.1 Data and Experimental Setup
We use TCGA patient cases which have available and valid survival data, pathology reports, tumor diagnostic slides, and tumor RNA-seq gene expression. For pathology reports, we use text extracted by kefeli2024tcga. For computational efficiency, we additionally rely on precomputed UNI2 embeddings of diagnostic slides by Chen et al. (chen2024towards). Given these requirements, we first filter TCGA cases by the availability of pathology report, diagnostic slide, then RNA-seq data. We download the extracted text from Mendeley (report-dataset), precomputed UNI2 embeddings of diagnostic slides from HuggingFace (uni2-dataset), and RNA-seq data using the Genomic Data Commons (GDC) API (heath2021nci). We ensure that the RNA-seq and diagnostic slides are derived from tumor samples.
2.2 Automated Report Summarization and Manual Correction
While the extracted pathology reports (kefeli2024tcga) are a rich source of information, they present multiple challenges. As the source reports are scanned documents of varying quality, the OCR extraction process can result in typos and loss of structured text. Additionally, the reports are often long with repeated information or information potentially irrelevant for cancer prognosis (e.g. incidental findings on histology).
To address these challenges, we experiment with having an LLM summarize the pathology reports to help focus relevant information, correct typos, and reduce report length. We choose Llama-3.1-8B-Instruct (grattafiori2024llama) as the summarization model for its strong instruction following capabilities. We use vLLM (kwon2023efficient) for inference with set seed, greedy decoding, temperature 0, and max tokens 1,024. Our prompt is presented in Table LABEL:tab:prompt.
As we experiment with model generated summaries of pathology reports, we test the effect of model hallucinations (huang2025survey) in the generated summaries and their downstream effects on survival modeling. To that end, we manually review model generated summaries for 40 randomly selected cases contained within a single test split. We randomly select these cases while preserving the overall prevalence of observed mortality. After review and manual correction (as needed) of the sampled summaries, we embed the corrected summaries using BioMistral and apply the survival model trained over BioMistral embeddings of summaries from the corresponding train split. We evaluate risk stratification over these sampled cases. Further details of our manual correction method are in Appendix A.
[…]
fig:inclusion Refer to caption
Figure 2: Eight thousand patient cases spanning 32 cancer types with survival data, pathology reports, diagnostic slides, and gene expression quantification.
[…]
@@ 2.2 Automated Report Summarization and Manual Correction
tab:unimodal Modality PCA Dimension 4 8 16 32 64 128 256 N/A Demo* - - - - - - - 0.630 Canc* - - - - - - - 0.737 Expr 0.626 0.650 0.702 0.742 0.749 0.753 0.750 - Hist 0.596 0.631 0.669 0.714 0.733 0.748 0.754 - Text 0.565 0.610 0.690 0.725 0.745 0.751 0.752 -
@@
Table 1: Foundation model derived embeddings predict cancer survival. Average cross-validated C-index of CoxPH models trained over unimodal embeddings across varying PCA reductions. Demo: demographics; Canc: cancer type; Expr: BulkRNABert embedded RNA-seq; Hist: UNI2 embedded histology; Text: BioMistral embedded summarized pathology reports. *Not reduced with PCA.
2.3 Foundation Model Embeddings
We use several foundation models in our experiments. For all FMs, we use the models as-is with no further training or fine-tuning. To prevent data leakage, we ensure that the FMs we use have not been trained over TCGA. Specifically, we experiment with UNI2-h (chen2024towards) for diagnostic slides, BulkRNABert (gelard2025bulkrnabert) or UCE (rosen2023universal) for gene expression data, and BioMistral-7B (labrak2024biomistral) or Mistral-7B-Instruct-v0.1 (jiang2023mistral) for pathology reports. Further details on these models is presented in Appendix B.
As survival is defined at the patient-level, we aggregate patient embeddings from multiple samples of the same modality into a single embedding per patient per modality. This is most relevant for RNA-seq and diagnostic slides as a patient may have multiple samples for these modalities. We use simple averaging of embeddings to derive our patient-level embeddings. Model embedding dimensionality shown in Table LABEL:tab:dimensions.
2.4 Unimodal and Multimodal Survival Modeling and Evaluation
Given patient-level, modality-specific embeddings, we adopt a simple pipeline for unimodal survival modeling. For a given train-test split, we z-score standardize all embeddings using per-feature mean and standard deviation derived from the training split. We next derive a Principal Component Analysis (PCA) (pearson1901liii) dimensionality reduction over the standardized embeddings from the training split and apply to all standardized embeddings. We use dimensionality reduction to prevent overfitting and improve computational tractability (see Table LABEL:tab:full-expr-hist and Appendix E for results and discussion on using full embeddings). As PCA is an efficient transformation that is insensitive to hyperparameter choices (as in t-SNE or UMAP), we specifically use the implementation of PCA from sklearn (pedregosa2011scikit) with a set seed. We experiment with varying PCA dimensions, doubling from 4 to 256. For demographic or cancer type modalities, features are one-hot encoded so we do not apply standardization or PCA transformation. We then fit a Cox Proportional Hazards (CoxPH) (cox1972regression) model to the PCA reduced embeddings (or the one-hot encoded features for demographic or cancer type modalities) of the train split. We use the implementation of CoxPH models from sksurv (polsterl2020scikit) with alpha = 0.1 for ridge regression penalty. Using the trained unimodal model, we predict risk scores for both the train and test splits for further modeling and evaluation.
We do late multimodal fusion using the predicted unimodal risk scores as input to the multimodal model. Given a set of modalities we aim to fuse, we concatenate the predicted risk scores from each modality as input features. We z-score standardize each feature using the mean and standard deviation derived over the training split. We fit a CoxPH model over the concatenated, standardized, unimodal risk scores. Such a model can be interpreted as modeling the unimodal risk (further discussed in Section 3.3). Using the trained multimodal model, we predict risk scores for the test split for evaluation. We repeat this multimodal fusion procedure over all combinations of demographic, cancer type, histology, expression, and text modalities. Notably, we do not consider alternate embedding methods as separate modalities; for example BioMistral embeddings of unsummarized and summarized reports are both text embeddings. Instead, we consider multimodal combinations with these alternate embedding strategies independently.
We evaluate the resulting models using 5-fold cross-validated concordance index (C-index) (harrell1982evaluating), mean cumulative/dynamic area under the receiver operating characteristic curve (mean AUCC,D, where larger is better) (lambert2016summary), integrated Brier score (IBS, where smaller is better) (brier1950verification), or risk stratified Kaplan-Meier (kaplan1958nonparametric) survival curves. Precise details on our evaluation method are provided in Appendix C.
3 Results
3.1 Multimodal data across 32 cancers
We assemble a dataset of 7,982 patient cases for our study (Figure LABEL:fig:inclusion). These cases span 32 of the 33 cancer types in TCGA. We select cases that have available and valid data for survival analysis, preextracted pathology reports by Kefeli and Tatonelli (kefeli2024tcga), precomputed tumor diagnostic slide embeddings by Chen et al. (chen2024towards), and tumor RNA-seq gene expression data.
For our experiments, we split the dataset into 5 stratified cross-validation folds that are shared across all experiments. Descriptive features of our dataset and splits are provided in Tables LABEL:tab:cohort and LABEL:tab:cancer-type. While we stratify by cancer type and observed mortality, we do not stratify by survival duration. On average, overall survival time is comparable across splits, however per-cancer type subsets contained within each split may have too few observed mortalities for survival modeling (Table LABEL:tab:cancer-type). See Appendix C for details on the limitations this presents for evaluating model performance within given cancer types.
[…]
tab:best-multimodal Modality PCA Dimension 4 8 16 32 64 128 256 N/A Canc*-Demo* - - - - - - - 0.747 Hist-Text 0.606 0.662 0.705 0.746 0.767 0.777 0.780 - Expr-Hist-Text 0.644 0.682 0.727 0.765 0.780 0.786 0.788 - Demo*-Expr-Hist-Text 0.680 0.717 0.748 0.776 0.788 0.794 0.795 - Canc*-Demo*-Expr-Hist-Text 0.749 0.752 0.760 0.777 0.788 0.793 0.793 -
Table 2: Mixing survival models of clinical features with FM-based survival models improves survival prediction. For the best combination of each number of modality combinations, average cross-validated C-index of multimodal CoxPH models trained over predicted risk scores from unimodal CoxPH models across varying PCA reductions. Demo: demographics; Canc: cancer type; Expr: BulkRNABert embedded RNA-seq; Hist: UNI2 embedded histology; Text: BioMistral embedded summarized pathology reports. *Not reduced with PCA.
3.2 Unimodal FM embeddings predict cancer survival
We fit small, linear CoxPH survival models over unimodal data and attain a peak mean cross-validated C-index of approximately 0.75 over single modalities (Table LABEL:tab:unimodal). For categorical demographic features or cancer type, we do no dimensionality reduction. As compared to the mixture of models over both clinical tabular features (Canc-Demo C-index = 0.747, Table LABEL:tab:best-multimodal), CoxPH models fit over demographic features or cancer type alone achieve slightly lower C-index (0.630 and 0.737, respectively).
For FM-derived embeddings (expression, histology, and text), we tested varying PCA sizes for dimensionality reduction. We observe a general trend of increasing C-index with increasing PCA dimensions, which plateaus around PCA to 256 dimensions. These findings are recapitulated using additional metrics of mean AUCC,D and IBS (Tables LABEL:tab:time-dependent-auroc and LABEL:tab:integrated-brier-score, respectively). Based on this observation, we report further results using PCA transformation of these embeddings to 256 dimensions.
With the same unimodal survival models, we evaluate their survival prediction performance across individual cancer types (Tables LABEL:tab:per-project-8 and LABEL:tab:per-project). We do not evaluate per-cancer survival based on cancer type as the input features would not differ within a cancer subset. We not only find that survival modeling performance of unimodal models varies across cancer types, we also observe that no single modality generally outperforms the other modalities. For example, among unimodal models, demographic-based modeling was most predictive of thyroid carcionma (THCA) survival (C-index = 0.885), while the same was true for text report-based modeling of endometrial carcinoma (UCEC) survival (C-index = 0.723).
3.3 Multimodal fusion improves cancer modeling and survival prediction